34 hours is taken by a worker to do job for which time allowed is 48 hours. TZS. 12000 per hour is his daily rate. Calculate the works cost of the job under the following methods of payment of wages:

(i) Halsey plan 50%

(ii) Halsey plan 33.3% &

(iii) Rowan plan.

Additional information: (i) Material cost TZS.600; (ii) Factory overhead 125% of wages


1
Expert's answer
2020-07-05T17:55:48-0400

1) Cost=(34*12000+0.5((48-34)*12000))*1.25+600=615600

2) Cost=(34*12000+0.33((48-34)*12000))*1.25+600=579900

3)Cost=(34*12000+ ((48-34)/48)*34*12000)*1.25+600=659350